OpenStudy (anonymous):

IMPORTANT: Need help with logarithms, am posting question with equations below. Medal to anyone who helps!

OpenStudy (anonymous):

Okay, so the question is, "Use change of base formula to find three significant digits." Here is the starting equation. \[\log_{x}y=\frac{ \log_{b}y }{\log_{b}x }\] The problem they want me to answer is \[\log_{2}6+\log_{3}5 \] Using the change of base formula, I plugged my info in and got \[\frac{ \log6 }{ \log2 }+\frac{ \log5 }{ \log3 }\] What I'm wanting to know is if, one, I set this up correctly, and two, if I can add the logs and if so how can I add the logs?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Okay, assuming you can compute or look up common logs (base 10), then yes, you have written this correctly and should be able to evaluate it now.

OpenStudy (anonymous):

What would be my next step?

OpenStudy (whpalmer4):

An interesting point is that it doesn't matter which log base you use for calculating the logs in the change of base formula: you've written it with the common log, but in fact you'll get the same answer if you do it with \(\log_{57}\), \(\ln\), or whatever you fancy :-)

OpenStudy (whpalmer4):

Bust out your calculator or go to a website which offers one and evaluate those logarithms. Or ask your friendly helper who memorized those values a few decades ago :-)

OpenStudy (anonymous):

Would I turn them into fractions?

OpenStudy (whpalmer4):

yes, but you have to find the values of the logs first

OpenStudy (whpalmer4):

\[\log 2 \approx 0.30103\] there's one to get you started :-)

OpenStudy (whpalmer4):

do you have a calculator handy with a logarithm button?

OpenStudy (whpalmer4):

\[\log 3 \approx 0.47721\] By the property of logs, we can use that to find \(\log 6\) \[\log a + \log b = \log(a*b)\]We know the values of \(\log 2\) and \(\log 3\), so \[\log 6 = \log(2*3) = \log 2 + \log 3 \approx 0.30103 + 0.47712 \approx 0.77815\]